OpenStudy (anonymous):

A skyscraper casts a shadow 200 feet long. If the angle of elevation of the sun is 45o , then the height of the skyscraper is ____ 170 feet 324 feet 200 feet not enough information

OpenStudy (anonymous):

Draw a triangle and use trignometry.

OpenStudy (goformit100):

not enough information

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

yeah here cos45 = hyp/200 so x here is 100(sqrt(2)) So the answer is A. 170 feet. Hope this helped.

OpenStudy (anonymous):

ty

jimthompson5910 (jim_thompson5910):

Boblovesmath, right idea, but wrong answer

jimthompson5910 (jim_thompson5910):

use tangent instead

jimthompson5910 (jim_thompson5910):

oh and it should be cos = adj/hyp

OpenStudy (anonymous):

oh yeah sorry, I found the hypotenuse, but I think the answer will be 170 feet as well.

OpenStudy (anonymous):

I am sorry the answer would be 324 feet, because hypotenuse always has the greatest magnitude in a triangle.

jimthompson5910 (jim_thompson5910):

tan = opp/adj tan(45) = x/200 solve for x